OpenStudy (anonymous):

The mass of mars is about 6.421 x 10^20 metric tons. The mass of the sun is about 1.998 x 10^27 metric tons. About how many times the mass of mars is the mass of the sun? GIve your answer in scientific notation.

OpenStudy (anonymous):

\[\frac{1.998 * 10^{27}}{6.421 * 10^{20}}=3.112*10^6 \]

OpenStudy (anonymous):

Thank you... I totally tried to solve that problem wrong!

OpenStudy (anonymous):

Your welcome.